OpenStudy (anonymous):

the distance (in mi) that a car travels in 45min; s=speed of the car (in m/hr)

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

that is the question that was in the book.

OpenStudy (anonymous):

that's now a question

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

\[\frac{3}{4}s\text{ }\text{miles} \]

OpenStudy (anonymous):

where did you get that from?

OpenStudy (anonymous):

there was no information

OpenStudy (anonymous):

r t = d s*(45/60) = D

OpenStudy (anonymous):

are you a robot? speak to me

OpenStudy (anonymous):

I understand the formula, but you have to talk

OpenStudy (anonymous):

rate * time = distance. rate = s miles per hour time has to be in hours 45 minutes = 45/60 hours Note: There are no numerical answers here.

OpenStudy (anonymous):

this was the question if I am correct -- the distance (in mi) that a car travels in 45min; s=speed of the car (in m/hr)

OpenStudy (anonymous):

that's not a question

OpenStudy (anonymous):

sorry and im not a robot >.<

OpenStudy (anonymous):

not you, I was talking to the guy who was answering, he doesn't talk, he just functions like a robot

OpenStudy (anonymous):

What I posted is the best solution given the initial conditions.

OpenStudy (anonymous):

what initial conditions?

OpenStudy (anonymous):

there were no conditions

OpenStudy (anonymous):

it's a statement

OpenStudy (anonymous):

The speed or rate is "s" mph. The time is 45 minutes or 0.75 hours. The distance traveled is 0.75 s .

OpenStudy (anonymous):

what distance are you talking about!

OpenStudy (anonymous):

#15

OpenStudy (anonymous):

angelita, copy what was before all those questions

OpenStudy (anonymous):

it's like one big question

OpenStudy (anonymous):

what distance are you talking about! The distance is ( 0.75 times s ) because "s" cannot be expressed as a number from the problem statement,

OpenStudy (anonymous):

\[\sum_{0}^{\infty} \sin(1/n^3)\]

OpenStudy (anonymous):

a series that could be represented in your function